# Physics-informed, boundary-constrained Gaussian process regression for the reconstruction of fluid flow fields

### v2

@@ 5.3.2 Discrete boundary constraints
Scalar BCGP kernel ${G}_{0}$} 415 0 830 $1.823\cdot 10^{-5}$
Scalar base kernel G{G} 344 142 830 $1.611\cdot 10^{-2}$
@@ 5.4 Flow around a NACA 0412 airfoil
In this section, we perform a GPR reconstruction of the velocity field of an incompressible fluid flow around the NACA 0412 airfoil at Reynolds number $\text{Re}=2.7\cdot 10^{3}$. We set slip conditions on the boundary of the airfoil, i.e. ${\boldsymbol{u}}({\boldsymbol{x}})\cdot{\boldsymbol{n}}({\boldsymbol{x}})=0$ for ${\boldsymbol{x}}\in{\Gamma_{0412}}$. The ground-truth data were generated using the AirfRANS dataset library [5] adapting for the slip boundary condition along the profile. A snapshot of the ground truth velocity field at $t=2$ s is depicted in Figure 9. The presented results are computed at this specific time step.
@@
Figure 9: Ground truth velocity field of an incompressible fluid flow around the NACA 0412 airfoil. The colormap represents the norm of the velocity field in m/s.
Along the same lines as in the previous section, we present in Figure 10 the GPR reconstructions of the velocity field using the BCGP-based approach and the discrete approach of Sect. 4.5 for the inclusion of the profile boundary condition. Since we consider a reconstruction problem of a ground-truth with high resolution, we set the computational domain as the rectangular box ${\Omega}=[-8,15]\times[-10,10]$ around the leading edge of the airfoil. The airfoil was fixed with chord length $c=1$ and the domain of the profile parameterization ${\boldsymbol{\gamma}}_{\text{NACA}}$} as ${\mathcal{P}}=[2.3,4]$, according to definition from Eq. (5.1).
Since this problem involves multi-scale dynamics of fluid motion, we define an additive anisotropic Gaussian kernel of the form:
${G}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\sum_{m=1}^{M}\sigma_{m}^{2}\exp\left[-\frac{1}{2}\left(\frac{(x_{1}-x_{1}^{\prime})^{2}+\alpha_{m}^{2}(x_{2}-x_{2}^{\prime})^{2}}{\ell_{m}^{2}}\right)\right],\quad{\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}\,.$ (5.5)
In our context, we consider this structure with two scales ($M=2$) as the base kernel for boundary-constraining. In particular, we set $\ell_{1}=0.64$ and $\alpha_{1}=8$ in the larger scale, while we fix $\ell_{2}=\ell_{1}/2^{6}$ and $\alpha_{2}=2$ to capture the motion in the smaller scale. We then set $\sigma_{2}=\sigma_{1}/2^{12}$ with $\sigma_{1}=0.4$ to adapt the kernel to the fluid energy decay, as explained in [34]. The computation of the associated BCGP kernel ${G}_{0}$} is done for each one of the choices of the measure $\nu$} from the Karhunen-Loève expansion over $\Gamma$, as in Corollaries 3.2 and 3.3.
The GPR observation design for each of the two approaches is done following a budget fixed by the Gram matrix size as ${N_{\text{Gram}}}=780$. In the BCGP-based case, we considered $N=390$ observations of velocity inside the domain, while in the discrete approach, we fixed $N=370$ positions for velocity information and ${N_{\Gamma}}=40$ discrete points along the boundary of the airfoil leading edge. These configurations along with the relative normal fit indicator from Eq. (4.15) are displayed in Table 3 for each approach. The nugget regularization parameter is fixed as $\eta=1\cdot 10^{-10}$. In Figure 11 a convergence study for the satisfaction of the profile boundary condition in terms of indicators from Eq. (4.14) and Eq. (4.15) with respect to the spectral accuracy of the kernel truncation is presented. Both choices of $\nu$} are taken into account for comparison. The numerical parameters for obtaining these estimates are $I=2000$ and $J+1=78$ (see Algorithms 1 and 2).
It is observed that under this specific configuration of kernel and profile geometry, the relative accuracy for the fulfillment of the slip boundary condition is better using BCGPs. Moreover, in this case, there is no GPR design which involves discrete observation of this condition in the posterior estimate.
(a) Ground truth
(b) GPR with BCGP kernel
(c) GPR with base kernel
(d) BCGP kernel
(e) Base kernel
Figure 10: Velocity field of an incompressible flow around a NACA 0412 airfoil leading edge. In (a) the ground truth is presented for comparison. In (b) the GPR estimate of the velocity field using the BCGP-approach for inclusion of the profile boundary condition is depicted using measure $\nu$} as in Corollary 3.2 (the estimate with $\nu$} as in Corollary 3.3 being visually equal). The counterpart using the discrete approach from Sect. 4.5 is presented in (c). The relative error fields of the GPR velocity estimates under these two approaches are presented in (d) and (e), respectively. The circumferences ($\circ$) are the GPR observation points of the velocity inside the computational domain $\Omega$}, whereas the stars ($\star$) are the observation points for the discrete design of the boundary condition on the airfoil. The colormap in (a) to (c) represents the norm of the velocity field in m/s, and the colormap in (d) and (e) corresponds to the norm of the relative error field.
GPR approach NN $N_{\Gamma}$} $N_{\text{Gram}}$} $\epsilon_{n}$
Base kernel G{G} 370 40 780 $5.76\cdot 10^{-3}$
BCGP kernel ${G}_{0}$} (pushforward $\nu$}) 390 0 780 $6.76\cdot 10^{-5}$
BCGP kernel ${G}_{0}$} (surface $\nu$}) 390 0 780 $8.02\cdot 10^{-5}$
Table 3: GPR design configurations and relative normal fit indicator $\epsilon_{n}$ from Eq. (4.15) for the GPR estimate of the velocity field around the leading edge of a NACA 0412 airfoil using multiple approaches for the inclusion of the slip boundary condition on the profile. The first approach considers the base kernel G{G} and ${N_{\Gamma}}=40$ discrete points for the observation of the slip boundary condition around the profile boundary. The other two approaches consider the associated BCGP kernel ${G}_{0}$} instead. Each one of them corresponds to a different choice of the measure $\nu$} from the Karhunen-Loève expansion, as described in Corollaries 3.2 and 3.3.
@@
(a) $\nu$} as the normalized pushforward measure ${{\boldsymbol{\gamma}}_{\text{NACA}}}_{*}{\eta}/\left\lvert{\mathcal{P}}\right\rvert$
@@
Figure 11: Evaluation of indicators of physical constraints around profile boundary with respect to spectral expansion convergence. The stream normal indicator of Eq. (4.14) is shown with stars ($\star$) and the velocity normal indicator $\epsilon_{n}$ of Eq. (4.15) is shown with circumferences ($\circ$). Both of these functions were computed by fixing the GPR design framework for the NACA 0412 leading edge with the same $N=390$ observations of velocity inside $\Omega$} and kernel hyperparameters. The horizontal axis is the negative logarithm of $\epsilon$ from Eq. (3.19) when varying the spectral bound $\delta$ in Algorithm 2. In (a) the measure $\nu$} of the Karhunen Loeve expansion is chosen as in Corollary 3.2, whereas in (b) the measure $\nu$} is chosen as in Corollary 3.3.

### v3

@@ 5.3.2 Discrete enforcement of the boundary conditions on the profile
RBF kernel G{G} 344 142 830 $1.611\cdot 10^{-2}$
BCGP kernel ${G}_{0}$} 415 0 830 $1.823\cdot 10^{-5}$
@@ 5.4 Flow around a NACA 0412 airfoil
We now turn to the GPR reconstruction of the velocity field of an incompressible fluid flow about a NACA 0412 airfoil at Reynolds number $\text{Re}=2.7\cdot 10^{3}$. The chord length of the profile ${\Gamma_{0412}}={\boldsymbol{\gamma}}({\mathcal{P}})$ is $c=1$ and the parameterization domain is ${\mathcal{P}}=[2.3,4]$, according to the definition of Eq. (5.1). We set slip conditions on the boundary of the airfoil, i.e. ${\boldsymbol{u}}({\boldsymbol{x}})\cdot{\boldsymbol{n}}({\boldsymbol{x}})=0$ for ${\boldsymbol{x}}\in{\Gamma_{0412}}$. The ground-truth data were generated using the AirfRANS dataset library [5] adapting for the slip boundary condition along the profile. A snapshot of the ground truth velocity field $\smash{{{\boldsymbol{u}}^{\dagger}}}$ is depicted in Figure 9. Since we consider a reconstruction problem with high resolution, the computational domain is set as the rectangular box ${\Omega}=[-0.08,0.15]\times[-0.1,0.1]$ around the leading edge of the airfoil. The presented results are computed using this specific stationary field.
(a) Leading edge
(b) Complete airfoil
@@
Figure 9: Stationary ground truth velocity field $\smash{{{\boldsymbol{u}}^{\dagger}}}$ of an incompressible fluid flow around the NACA 0412 airfoil. The colormap represents the norm of the velocity field in m/s.
Our objective is to compare different choices of GP priors used in the reconstruction formula (4.11) through the definition of their respective kernels. Multiple physical conditions can be included at this stage. We use the baseline kernel structure introduced in Eq. (4.9). The parameter $M$ stands for the number of fine scales used to capture the energy decay of the fluid dynamics. The following three cases are investigated:
  1. 1.
    Unique coarse scale using a RBF: an exponential kernel, as defined by Eq. (4.8) (or Eq. (4.9) with $M=0$);
  2. 2.
    Multiple scales using RBF (M-RBF): an exponential kernel informed with energy decay, as defined by Eq. (4.9) with $M=3$;
  3. 3.
    Physics-informed BCGP (PI-RBF): a multiple-scale exponential kernel informed about the energy decay and the profile boundary condition from Eq. (4.7), through the BCGP procedure, where the base kernel is defined as in Eq. (4.9) with $M=3$. The number of integration intervals is $I=300$ and the spectral accuracy bound $\delta$ is set to $\delta=10^{-12}$, which yields 112 spectral modes (see Algorithms 1 and 2). The spectral measure $\nu$} is chosen as in Corollary 3.3. A comparison with respect to the measure from Corollary 3.2 is presented later in Figure 14.
The hyperparameters $\sigma_{0}$ and $\ell_{0}$ for the standard deviation and correlation length, respectively, are estimated using a cross-validation methodology as depicted below in Sect. 5.4.1. The regularization parameter $\eta$ is set at $\eta=\smash{10^{-8}}$ for the RBF kernel and at $\eta=\smash{10^{-10}}$ for the M-RBF and PI-RBF kernels to ensure that the Gram matrices are invertible.
The training set $\boldsymbol{X}$} gathers $N=184$ velocity measurement positions which are selected randomly. No observation positions are considered near or at the boundary of the NACA airfoil, though. The test set $\smash{{{\boldsymbol{X}}_{\text{test}}}}=\smash{\{{{\boldsymbol{x}}^{k}_{{\Omega}}}\}_{k=1}^{N_{\text{test}}}}$ gathers ${N_{\text{test}}}=3439$ positions inside the computational domain. These positions have a higher concentration near the profile for a better resolution of the reconstructed fields. Note that the same training and test sets are used in order to compare the three GPR reconstruction approaches described above. The measure of goodness of fit over the test set is the root mean squared error $\smash{\epsilon_{\text{RMSE}}}$ defined as:
$\epsilon_{\text{RMSE}}({{\boldsymbol{X}}_{\text{test}}};{{\boldsymbol{u}}^{\star}})=\sqrt{\frac{1}{{N_{\text{test}}}}\sum_{k=1}^{{N_{\text{test}}}}\left\|{{\boldsymbol{u}}^{\star}}({{\boldsymbol{x}}^{k}_{{\Omega}}})-{{\boldsymbol{u}}^{\dagger}}({{\boldsymbol{x}}^{k}_{{\Omega}}})\right\|^{2}}\,$, (5.6)
where $\smash{{{\boldsymbol{u}}^{\dagger}}}$ is the velocity field ground truth, and $\smash{{{\boldsymbol{u}}^{\star}}}$ is the velocity field from the reconstruction formula (4.11) associated to a specific kernel definition.
In order to quantify the uncertainty associated to these different reconstructions, we introduce the UQ coverage indicators $\smash{p_{1}}$ and $\smash{p_{2}}$ defined as the posterior coverage percentages for the horizontal and vertical velocity components, respectively. These quantities are computed as the ratios of test points lying inside their corresponding posterior $95\%$ confidence intervals. That is, given the set $\smash{{{\boldsymbol{X}}_{\text{test}}}}$ of $N_{\text{test}}$} locations within the computational domain and a reconstruction formula $\boldsymbol{u}$$^{\star}$}, the indicators are defined by $\smash{p_{i}({{\boldsymbol{X}}_{\text{test}}};{{\boldsymbol{u}}^{\star}})}=\smash{N_{i}/{N_{\text{test}}}}$ with:
$N_{i}=\#\left\{{\boldsymbol{x}}\in{{\boldsymbol{X}}_{\text{test}}}\,:\,\left\lvert\smash{u^{\star}_{i}({\boldsymbol{x}})-u_{i}^{\dagger}({\boldsymbol{x}})}\right\rvert\leq 1.96\times\sigma_{i}({\boldsymbol{x}})\right\}\,$, (5.7)
for $i=1,2$. Here $\sigma_{i}({\boldsymbol{x}})=\smash{\sqrt{\Sigma_{ii}^{\star}({\boldsymbol{x}},{\boldsymbol{x}})}}$ is the posterior standard deviation of the $i$-th velocity component at ${\boldsymbol{x}}\in{\Omega}$, for $\smash{\Sigma_{ii}^{\star}({\boldsymbol{x}},{\boldsymbol{x}})}$ being the $i$-th element of the diagonal of the matrix-valued covariance ${\boldsymbol{\Sigma}^{\star}}({\boldsymbol{x}},{\boldsymbol{x}})$ from Eq. (4.13).
5.4.1 Kernel hyperparameter estimation with UQ-based cross-validation
In order to estimate the remaining hyperparameters, we use the training set to perform cross-validation with a loss function defined in terms of the UQ coverage indicators from Eq. (5.7). To do so, we split the training set $\boldsymbol{X}$} into $K$ disjoint folds of equal size: ${\boldsymbol{X}}=\smash{{\boldsymbol{X}}^{(1)}\cup\cdots\cup{\boldsymbol{X}}^{(K)}}$. Then we define the individual loss function $\mathcal{L}$} given a validation fold $\smash{{\boldsymbol{X}}^{(k)}}$ and a selection of hyperparameters $\smash{\sigma_{0},\ell_{0}}$ (which are associated to a reconstruction formula $\boldsymbol{u}$$^{\star}$}):
${\mathcal{L}}({\boldsymbol{X}}^{(k)};\sigma_{0},\ell_{0})=\frac{1}{2}\left[\left(p_{1}({\boldsymbol{X}}^{(k)};{{\boldsymbol{u}}^{\star}})-0.95\right)^{2}+\left(p_{2}({\boldsymbol{X}}^{(k)};{{\boldsymbol{u}}^{\star}})-0.95\right)^{2}\right]\,.$ (5.8)
We use (5.7) for the computation of $\smash{p_{1}}$ and $\smash{p_{2}}$, replacing the test set $\smash{{{\boldsymbol{X}}_{\text{test}}}}$ by each validation fold ${\boldsymbol{X}}^{(k)}$. Finally, we perform grid search to minimize the mean loss function:
${\mathcal{L}}_{K}(\sigma_{0},\ell_{0})=\frac{1}{K}\sum_{k=1}^{K}{\mathcal{L}}({\boldsymbol{X}}^{(k)};\sigma_{0},\ell_{0})\,.$ (5.9)
In our experiment, we use $K=4$ folds to get estimates of the hyperparameters $\sigma_{0}$ and $\ell_{0}$ for the three investigated kernels by minimizing over the grids displayed in Figure 10. The logarithm of the mean loss function over these grids is depicted for each kernel. The minimal nodes are selected for computing the reconstructions in the next sections.
(a) RBF kernel ($M=0$)
(b) M-RBF kernel ($M=3$)
(c) PI-RBF kernel ($M=3$)
[…]
5.4.2 Flow reconstruction with uncertainty quantification
Figure 11(a)-(c) display the reconstructed velocity fields $\boldsymbol{u}$$^{\star}$} within the computational domain $\Omega$} using each one of the three kernels (RBF, M-RBF, and PI-RBF). The corresponding residual fields $\smash{{{\boldsymbol{u}}^{\star}}-{{\boldsymbol{u}}^{\dagger}}}$ are depicted in Figure 11(d)-(f). These reconstructions are performed using the same training set $\boldsymbol{X}$}, stressing again that no velocity measurements are set at the boundary of the profile at all. A summary of the regression configurations as well as the fit evaluation indicators are gathered in Table 3. The last rows correspond to the UQ coverage indicators computed on the test set ${\boldsymbol{X}}_{\text{test}}$}. We notice that the reconstruction improves from RBF to M-RBF by the inclusion of larger scales and a proper choice of hyperparemeters. However, there is still no means of getting information about the profile boundary with either of these two approaches. The boundary condition fit improves significantly with the inclusion of the BCGP procedure. In this way, even if there is no discrete observation on the profile boundary for the computation of the posterior estimates, the PI-RBF kernel is able to recover this information. The numerical indicator $\epsilon_{n}$ of Eq. (4.18) quantifying the goodness of fit of the boundary condition is significantly decreased (by three orders of magnitude) compared to the other kernels. The choice of a PI-RBF kernel improves the reconstructed velocity fields at the leading edge of the profile as well.
(a) RBF kernel ($M=0$)
(b) M-RBF kernel ($M=3$)
(c) PI-RBF kernel ($M=3$)
(d) RBF kernel ($M=0$)
(e) M-RBF kernel ($M=3$)
(f) PI-RBF kernel ($M=3$)
Figure 11: Reconstruction of the velocity field of an incompressible flow at the leading edge of a NACA 0412 airfoil using GPs with different kernels. In the first row from (a) to (c), the norm of the reconstructed velocity field $\boldsymbol{u}$$^{\star}$} is presented in m/s. In the second row from (d) to (f), the norm of the residual field $\smash{{{\boldsymbol{u}}^{\star}}-{{\boldsymbol{u}}^{\dagger}}}$ is presented in m/s. The three choices of GP kernels are: one-scale RBF in (a) and (d); multiple-scale M-RBF accounting for energy decay in (b) and (e); and physics-informed multiple-scale PI-RBF (informed about the profile boundary condition through the BCGP procedure) in (c) and (f). The black dots ($\bullet$) are the velocity observation positions corresponding to the training set within the computational domain $\Omega$}.
GP kernel approach RBF M-RBF PI-RBF
Kernel hyperparameters $M=0$ $\sigma_{0}=0.01$ $\ell_{0}=0.1$ $M=3$ $\sigma_{0}=0.05$ $\ell_{0}=1.0$ $M=3$ $\sigma_{0}=0.10$ $\ell_{0}=1.0$
Regularization parameter $\eta$ $^{-8} 1010$$^{-10}$ $10^{-10}$
$\epsilon_{n}$ from Eq. (4.18) $7.9\cdot 10^{-2}$ $4.8\cdot 10^{-2}$ $9.3\cdot 10^{-6}$
$\epsilon_{\text{RMSE}}$ from Eq. (5.6) $4.0\cdot 10^{-3}$ $2.7\cdot 10^{-3}$ $7.6\cdot 10^{-4}$
UQ coverage indicators on ${\boldsymbol{X}}_{\text{test}}$} $p_{1}=81.7\%$ $p_{2}=78.2\%$ $p_{1}=78.2\%$ $p_{2}=84.5\%$ $p_{1}=96.7\%$ $p_{2}=95.3\%$
Table 3: Configuration and evaluation for the reconstruction of the velocity field of an incompressible flow at the leading edge of a NACA 0412 airfoil using GPs with different kernels. The three choices of GP kernels are: one-scale RBF; multiple-scale M-RBF accounting for energy decay; and physics-informed multiple-scale PI-RBF (informed about the profile boundary condition through the BCGP procedure). The kernel hyperparameters are set through cross-validation using the training set; see Sect. 5.4.1. The UQ coverage indicators of each velocity component are computed using the test set and the corresponding velocity reconstruction formula of each configuration.
In order to quantify the uncertainty associated to the velocity field estimates, the (logarithmic) root-mean-square standard deviations ${\boldsymbol{x}}\mapsto\smash{\log_{10}\sqrt{\sigma_{1}^{2}({\boldsymbol{x}})+\sigma_{2}^{2}({\boldsymbol{x}})}}$ are displayed in Figure 12. It can be observed that the inclusion of boundary information reduces the uncertainty in the part of the computational domain surrounding the airfoil. Moreover, the posterior confidence intervals of each velocity component are displayed in Figure 13. The inclusion of multiple energy scales and the profile boundary condition significantly improve the fit on both components of velocity and yield a consistent calibration of the confidence intervals. As explained in the previous sections, the BCGP procedure enforces the information from Eq. (4.7) on all points of the profile boundary. The tangential component of velocity are left to be learned from data in the posterior distribution only. The posterior variance on this boundary is therefore not null, but significantly consistent with the $95\%$ confidence precision.
(a) RBF kernel ($M=0$)
(b) M-RBF kernel ($M=3$)
(c) PI-RBF kernel ($M=3$)
Figure 12: UQ of the reconstructed velocity field of an incompressible flow at the leading edge of a NACA 0412 airfoil using GPs with different kernels. The depicted fields are the base 10 logarithm of the posterior total standard deviation. The three choices of GP kernels are: one-scale RBF; multiple-scale M-RBF accounting for energy decay; and physics-informed multiple-scale PI-RBF (informed about the profile boundary condition through the BCGP procedure). The black dots ($\bullet$) are the velocity observation positions corresponding to the training set within the computational domain $\Omega$}.
(a) RBF kernel ($M=0$)
(b) M-RBF kernel ($M=3$)
(c) PI-RBF kernel ($M=3$)
(d) RBF kernel ($M=0$)
(e) M-RBF kernel ($M=3$)
(f) PI-RBF kernel ($M=3$)
Figure 13: Estimates and confidence intervals of the horizontal (from (a) to (c)) and vertical (from (d) to (f)) components of the velocity field at the leading edge of a NACA 0412 airfoil. The horizontal coordinates are the parameterization domain ${\mathcal{P}}=[2.3,4]$. The three choices of GP kernels are: one-scale RBF; multiple-scale M-RBF accounting for energy decay; and physics-informed multiple-scale PI-RBF (informed about the profile boundary condition through the BCGP procedure). The black dashed line represents the ground truth of the velocity field components. The continuous blue line represents the posterior mean components. The gray-shaded areas are the posterior $95\%$ confidence intervals.
Finally, we compare in Figure 14 the satisfaction of the profile boundary condition when using the PI-RBF kernel for two choices of the spectral measure $\nu$} (see Sect. 3.2). The indicators defined in Eq. (4.17) and Eq. (4.18) are presented for both choices of measure. The convergence is presented on the grid $\left\{10^{-n}\,:\,n=9,\,10,\,\ldots,14\right\}$ corresponding to the spectral accuracy bound $\delta$ of the kernel truncation (see Algorithms 1 and 2). The discretization of the integral operator was set at $I=300$ intervals. The resulting number of spectral modes ranged from 23 at $\delta=10^{-9}$ to 200 at $\delta=10^{-14}$.
@@ 5.4.2 Flow reconstruction with uncertainty quantification
(a) $\nu$} as the normalized pushforward measure ${\boldsymbol{\gamma}}_{*}{\eta}/\left\lvert{\mathcal{P}}\right\rvert$
@@
Figure 14: Evaluation of indicators of physical constraints around profile boundary with respect to spectral expansion convergence. The stream normal indicator of Eq. (4.17) is shown with stars ($\star$) and the velocity normal indicator $\epsilon_{n}$ of Eq. (4.18) is shown with circles ($\circ$). Both of these functions were computed by fixing the GPR design framework for the NACA 0412 leading edge with the same training set and kernel hyperparameters. The horizontal axis is the negative logarithm of $\epsilon$ from Eq. (3.19) when varying the spectral bound $\delta$ in Algorithm 2. In (a) the measure $\nu$} of the Karhunen-Loève expansion is chosen as in Corollary 3.2, whereas in (b) the measure $\nu$} is chosen as in Corollary 3.3.